# CRUD-RAG: A Comprehensive Chinese Benchmark for Retrieval-Augmented Generation of Large Language Models

### v2

@@ 3.1. News Collection
We argue that the latest news data is the most suitable choice to create an RAG evaluation dataset. Unlike other types of data, such as encyclopedias, questions, or conversations, the latest news data prevent the model from generating answers directly from its own knowledge base. Instead, the model has to use external documents to generate content. This allows us to evaluate the performance of the overall RAG pipeline, rather than just the model’s generation capabilities. Moreover, the latest news data offer rich and diverse topics and content, which can test the model’s performance and adaptability in various domains and situations.
Therefore, we select news as the base of our datasets. To ensure the authenticity and currency of the datasets, we collected nearly $300,000$ of historical news articles from major Chinese news websites published after July 2023, which were not exposed to the LLMs during the training phase. We remove duplicate news documents from $300,000$ news articles and filter out those that were too long or too short. We ended up with more than $80,000$ news articles. Based on the news corpus we collected, we constructed our datasets for three tasks, namely open-domain multi-document summarization, text continuation, and question answering.
@@ 3.2. Open-domain Multi-document Summarization: RAG Application in ”Delete”
However, constructing such a dataset is extremely challenging. First, news articles retrieved based on events may not be fully relevant, requiring manual filtration to identify the correct and pertinent documents. Then, when generating summaries from these documents, it is essential to eliminate a significant amount of redundant information, retaining only the most important content. These tasks require manual annotation, which consumes substantial time and financial resources, and often result in too much superfluous information.
@@ 3.3. Text continuation: RAG Application in ”Create”
To construct the continuation task dataset, we follow the same method as the summary task dataset. Figure 4 shows the construction process of text continuation. Specifically, we select a news article from a high-quality corpus, and use a specialized Chinese word segmentation tool, to split it into sentences. Then, we divide the article into two equal parts: the first half serves as the input and the second half as the output of the continuation dataset. We expect the model to use RAG technology to retrieve relevant information from the document library and generate a continuation that is coherent, informative, and consistent with the input and output.
To ensure that the retrieval database covers the real continuation text, we use Baidu search engine to find external documents and add them to the database. The continuation text differs from the event text in that it consists of multiple sentences. Therefore, we split the continuation text into paragraphs by sentences and retrieve relevant documents for each paragraph using search engine. This way, we guarantee that the retrieval database contains most of the information to reconstruct the continuation text.
@@ 3.4. Question Answering: RAG Application in ”Read”
Another application scenario of RAG is to use external knowledge bases to enhance the question-answering capabilities of large language models, which can be applied to various knowledge-intensive tasks. Currently, there are many evaluation benchmarks to measure the performance of RAG in this scenario, and multiple question answering datasets have been created.
However, the existing question answering datasets also have some limitations. On the one hand, some datasets (such as NQ and WEBQA) are outdated, and may have been covered by large language models in the pre-training stage, which reduces the advantage of RAG systems. On the other hand, some datasets (such as RGB) only contain some factual questions, which can be directly extracted from the retrieved texts, without requiring complex reasoning over multiple texts, which poses less challenge to RAG systems. The most recent LLMs capture enough knowledge to rival human performance across a wide variety of question answering benchmarks (Bubeck et al., 2023).
@@
For the single-document question answering task, we follow the dataset construction process of the previous RGB benchmark (Chen et al., 2023b). We first select news articles from our collected high-quality corpus. Then we use prompts to make ChatGPT generate questions and answers for each article. For example, for a report on ”The 2023 Nobel Prize”, ChatGPT will generate the question ”Who was awarded the 2023 Nobel Prize for Physiology and Medicine?” and provide key information for answering it.
For the multi-document question answering task, constructing a reasoning question that requires the synthesis of multiple documents is not trivial. Simply using a prompt to force ChatGPT to generate the question is ineffective, because creating such a multi-document QA dataset is a complex reasoning task in itself. Therefore, we adopt Chain-of-Thought (CoT) technology (Wei et al., 2022) to enhance ChatGPT. We guide the model to build the dataset gradually through multiple reasoning steps. Figure 5 illustrates our specific process for building a two-document question answering dataset using ChatGPT and CoT technology. We will explain it in detail:
@@
    Generate question based on different focus points, which should require integrating information from multiple sources to answer.
@@ 3.6. Evaluation Method
Evaluating the performance of RAG systems requires choosing appropriate evaluation metrics. We considered the previous evaluation metrics for text generation, ROUGE and BLEU, which are both based on word overlap. ROUGE mainly counts the recall rate on n-gram, while BLEU mainly counts the precision rate on n-gram. However, BLEU and ROUGE are word overlap-based accuracy metrics that depend on the overall expression of the text, and do not capture the accuracy of the particular key information in the text. Therefore, they may not reflect the factual consistency of a text well, especially for long texts. To alleviate this issue, recent work (Wang et al., 2020; Scialom et al., 2021; Durmus et al., 2020) has proposed new evaluation metrics for abstractive summarization evaluation. These metrics are based on the intuition that if you ask questions about the summary and the original document, you will get a similar answer if the summary realistically matches the original document. They evaluate the accuracy of each local piece of key information in the summary.
@@
We also consider question answering-based metrics to evaluate the factual accuracy of generation. In this paper, we examine QuestEval (Scialom et al., 2021), a metric that improves the correlation with human judgments over previous metrics in their extensive experiments. QuestEval evaluates the factual consistency between the generated text and the source document, which is mainly used for text summarization tasks. Therefore, it does not require any ground truth reference. However, for RAG systems, the retrieved texts may be irrelevant or incorrect, so consistency with them is not a valid criterion. Instead, we use this metric to measure how well the generated text matches the ground-truth reference. We call this metric RAGQuestEval. We will explain this metric in detail.
@@ 4.1. Experimental Settings
  • Top-k: This is the number of text chunks that the RAG system retrieves for each query, which serve as the context part of the LLMs prompts. The top-k influences the size of the context that the model receives.
@@
Table 3. The experimental results for evaluating different chunk size in our benchmark. We use two types of evaluation metrics: The overall semantic similarity metrics(bleu, rouge-L, and bertScore) and the key information metric(RAGQuestEval).
@@ 4.2. Analyzing the Impact of Chunk Size on RAG Performance in Different Tasks
Chunking is the process of dividing a document into chunks of a fixed length, and then converting each chunk into a vector and storing it in an index. This creates an external knowledge index. Chunk size is a crucial parameter that varies depending on the corpus characteristics. If the chunk is too small or too large, it can reduce the search accuracy or omit important content. Hence, finding the optimal chunk size is vital for ensuring the search accuracy and relevance, and enabling the LLMs to generate appropriate responses. Our experiments reveal that different RAG tasks correspond to different optimal chunk sizes.
@@
Text Continuation: The experimental results in Table 3 demonstrate that larger chunk size can significantly improve the overall semantic similarity measures (bleu, rouge-L, bertScore). Besides, the RAGQuestEval metrics, which reflect the precision and recall rate of key information, follow a consistent pattern. This indicates that larger blocks preserve the original document’s structure, which is crucial for creative tasks such as text continuation. Smaller chunks, on the other hand, result in fragmented and semantically incoherent content, which impairs the ability of large models to understand and generate engaging content.
@@
Open-Domain Multi-Document Summarization: We observe some intriguing patterns in the experimental results. Firstly, we discover that larger chunk size not only substantially increase the length of the generated text, but also cause a notable drop in the bleu score, while the rouge-L and bertScore remain almost unchanged. This implies that larger chunks can preserve more original text information, but also introduce some semantic redundancy. Secondly, for the RAGQuestEval metric that evaluates key information, we found that a larger chunk size considerably enhances the recall of key information, but also lowers the precision of key information.
@@
We hypothesize that this is because larger blocks enable the retrieval of more relevant content without breaking the article structure, thus improving the recall of key information. However, larger blocks also make the summarization task more challenging, as more fine-grained selection is required from the more relevant information, leading to lower precision of key information.generated text, which may not be a good thing for the summary.
@@
Question Answering: For single-document QA, larger blocks will reduce both recall and precision score of key information. The task only requires extracting information from a sub-paragraph of a single document, and the answer may be in a specific sentence. Therefore, smaller chunks are more suitable, as excessive content will make the extraction harder for the model.
@@
For multi-document QA, the results are different from those of single-document QA. Larger blocks can significantly improve the recall and precision of key information, as well as the semantic similarity of the generated and reference answers. This is because larger blocks retain the original structure of the article, which is crucial for reasoning understanding tasks, and fragmented information is not conducive to reasoning.
@@
Hallucination Modification: For the hallucination modification task, the results are similar to those of the single-document QA task. Smaller blocks can significantly improve the semantic similarity metrics, such as BLEU score. This indicates that in the hallucination dataset created by UHGEval, the hallucination information often pertains to only one sentence, which is a mistake at the word or entity level, and does not require the comprehension of long text. Hence, there is no need to understand the whole document, only the relevant portions can be retrieved and modified.
@@ 4.3. Analyzing the Impact of Chunk Overlap on RAG Performance in Different Tasks
Chunk overlap is the number of tokens that two adjacent chunks share. To keep the text semantics coherent, adjacent chunks have some overlap. Chunk overlap determines the size of this overlap. This splitting method meets the maximum length limit of LLMs and maintains the semantic connection between adjacent chunks. Suitable chunk size and overlap can enhance the fluency and coherence of large language models for long texts. We will show how the chunk overlap rate affects the system performance for different tasks in the Table 4.
Text Continuation: With an increase in chunk overlap, we observe significant enhancements in the metrics that evaluate the alignment of generated text with a reference answer (BLEU, ROUGE, and BERTScore). The RAGQuestEval metric, which evaluates the accuracy and completeness of important information, also improves. These results indicate that a greater chunk overlap is beneficial for preserving the flow of ideas in the text, which is essential for tasks that require generating new, creative content.
@@
Question Answering: The consistency of meanings across diffierent chunks is especially important in question answering tasks. Data shows that as chunk overlap increases, the metrics of precision and recall of key information improve significantly in the single-document question answering task. A similar improvement is evident for two-document question answering. However, for three-document question answering, the enhancements are more modest. This might be because these tasks require a wealth of context, and a large chunk overlap could cut down on the context available.
@@ 4.4. Analyzing the Impact of Retriever on RAG Performance in Different Tasks
Text Continuation: As Table 5 displays, dense retriever outperforms BM25. Compared to the keyword-based algorithm, the modern vector search can capture the semantic and contextual information of words, and greatly enhance the retrieval performance. Thus, in the RAG pipeline, the generation results using dense retriever are also better. However, the RAG system using BM25 also performs well. In terms of the precision of key information, BM25 even exceeds Dense retriever. This suggests that in the continuation task, which is a creative task, BM25 can retrieve content that is highly relevant to the user’s intention, but may overlook some details.
Open-Domain Multi-Document Summarization: On the overall semantic similarity metric, BM25 outperforms dense retriever. On the QuestEval metric, BM25 surpasses dense retriever in terms of key information precision, but slightly trails behind in key information recall. If the retrieved content contains a lot of irrelevant information, the model-generated summary may have errors or redundancies. BM25 retrieved content usually matches the user’s intention better, but sometimes may miss some important information. Therefore, BM25 is slightly weaker than dense retriever in key information recall, but excels in key information accuracy. Besides, hybrid retrieval algorithms presumably combine the advantages of both, and the RAG system generates content with suitable precision and recall.
Question Answering: In question answering, we find that dense retriever has a more obvious advantage over BM25, when dealing with reasoning questions that require synthesizing multiple documents. In question answering tasks that require considering three documents, Dense retriever not only surpasses BM25 in all the overall semantic similarity metrics, but also achieves a significant improvement in key information precision and recall. This indicates that question answering retrieval is more difficult than text continuation and other tasks, especially reasoning question answering, which requires a higher level of semantic understanding, and simple keyword retrieval algorithms may not be sufficient. We also found that the Hybrid+Rerank algorithm, which combines and re-ranks the results of both algorithms, improves on all evaluation metrics. This suggests that this is a better retrieval algorithm for question answering tasks.
Hallucination Modification: Consistent with the conclusion of summarization, BM25 retriever performs slightly better than dense retriever, and the hybrid retrieval algorithm presumably combines the advantages of both. For RAG tasks such as hallucination modification, which require precise retrieval of highly relevant content, BM25 shows greater advantages. Moreover, BM25 requires less computational resource then dense retriever. This indicates that different RAG tasks require different retrieval algorithms

### v3

@@ 3.1. News Collection
We argue that the latest news data is the most suitable choice for creating an RAG evaluation dataset. Unlike other types of data, such as encyclopedias, questions, or conversations, the latest news minimizes the possibility that the model has been exposed to similar content during training. This dependency on external retrieval mechanisms allows for a comprehensive evaluation of the entire RAG process, not just the model’s generation ability. Additionally, news data is easy to collect, enabling us to maintain dataset timeliness. When the existing dataset loses its timeliness, we can quickly gather the latest news to rebuild a more challenging dataset. Moreover, the latest news data offer rich and diverse topics and content, which can test the model’s performance and adaptability in various domains and situations.
Therefore, we select news as the base of our datasets. To ensure the authenticity and currency of the datasets, we collected nearly $300,000$ of historical news articles from major Chinese news websites published after July 2023, which were not exposed to the LLMs during the training phase. We remove duplicate news documents from $300,000$ news articles and filter out those that are too long or too short. We ended up with more than $80,000$ news articles. Based on the news corpus we collected, we constructed our datasets for three tasks, namely open-domain multi-document summarization, text continuation, and question answering.
@@ 3.2. Open-domain Multi-document Summarization: RAG Application in ”Delete”
However, constructing such a dataset is extremely challenging. First, news articles retrieved based on events may not be fully relevant, requiring manual filtration to identify the correct and pertinent documents. Then, when generating summaries from these documents, it is essential to eliminate a significant amount of redundant information, retaining only the most important content. These tasks require manual annotation, which consumes substantial time and financial resources, and often results in too much redundant information.
@@ 3.3. Text continuation: RAG Application in ”Create”
To construct the continuation task dataset, we follow the same method as the summary task dataset. Figure 4 shows the construction process of text continuation. Specifically, we select a news article from a high-quality corpus and use a specialized Chinese word segmentation tool, to split it into sentences. Then, we divide the article into two equal parts: the first half serves as the input, and the second half as the output of the continuation dataset. We expect the model to use RAG technology to retrieve relevant information from the document library and generate a continuation that is coherent, informative, and consistent with the input and output.
To ensure that the retrieval database covers the real continuation text, we use the Baidu search engine to find external documents and add them to the database. The continuation text differs from the event text in that it consists of multiple sentences. Therefore, we split the continuation text into paragraphs by sentences and retrieve relevant documents for each paragraph using the search engine. This way, we guarantee that the retrieval database contains most of the information to reconstruct the continuation text.
@@ 3.4. Question Answering: RAG Application in ”Read”
Another application scenario of RAG is to use external knowledge bases to enhance the question-answering capabilities of LLMs, which can be applied to various knowledge-intensive tasks. Currently, there are many evaluation benchmarks to measure the performance of RAG in this scenario, and multiple question answering datasets have been created.
However, the existing question answering datasets also have some limitations. On the one hand, some datasets (such as NQ and WEBQA) are outdated, and may have been covered by LLMs in the pre-training stage, which reduces the advantage of RAG systems. On the other hand, some datasets (such as RGB) only contain some factual questions, which can be directly extracted from the retrieved texts, without requiring complex reasoning over multiple texts, which poses less challenge to RAG systems. The most recent LLMs capture enough knowledge to rival human performance across a wide variety of question answering benchmarks (Bubeck et al., 2023).
@@
For the single-document question answering task, we follow the dataset construction process of the previous RGB benchmark (Chen et al., 2023b). We first select news articles from our collected high-quality corpus. Then we use prompts to make GPT-4 generate questions and answers for each article. For example, for a report on ”The 2023 Nobel Prize”, GPT-4 will generate the question ”Who was awarded the 2023 Nobel Prize for Physiology and Medicine?” and provide key information for answering it.
For the multi-document question answering task, constructing a reasoning question that requires the synthesis of multiple documents is not trivial. Simply using a prompt to force GPT-4 to generate the question is ineffective, because creating such a multi-document QA dataset is a complex reasoning task in itself. Therefore, we adopt Chain-of-Thought (CoT) technology (Wei et al., 2022) to enhance GPT-4. We guide the model to build the dataset gradually through multiple reasoning steps. Figure 5 illustrates our specific process for building a two-document question answering dataset using GPT-4 and CoT technology. We will explain it in detail:
@@
    Generate the question based on different focus points, which should require integrating information from multiple sources to answer.
@@
To further ensure our dataset’s quality, we employed a manual refinement process for the data generated by GPT-4. Our annotation team comprises three native Chinese speakers, each with at least a bachelor’s degree. The annotation process is as follows:
  1. (1)
    The annotator evaluates the quality of the automatically generated query and chooses one of the following two options:
    • Reasonable: Conforms to natural language usage.
    • Needs refinement: Has issues with naturalness, accuracy, or grammar.
  2. (2)
    If ”Reasonable” is selected, no further action is taken. If ”Needs refinement” is chosen, the annotator manually improves the query’s naturalness and accuracy.
In addition to their standard salary, annotators receive an extra 1 RMB per query evaluated or refined. The average annotation time per query is approximately 20 seconds. To ensure annotation quality, we randomly inspected 5% of the annotated data.
Given the substantial cost of manual annotation and the large size of our dataset, we initially polished one-fifth of our dataset manually. We will continuously monitor dataset quality across various social media platforms and refine it manually as needed.
Notably, only 5.8% of queries required refinement, indicating that the queries generated by GPT-4 are generally of high quality. This validates the effectiveness of using GPT-4 for initial data generation and underscores our commitment to ensuring dataset quality.”
@@ 3.6. Evaluation Method
Evaluating the performance of RAG systems requires choosing appropriate evaluation metrics. We considered the previous evaluation metrics for text generation, ROUGE and BLEU, which are both based on word overlap. ROUGE mainly counts the recall rate on the n-gram, while BLEU mainly counts the precision rate on the n-gram. However, BLEU and ROUGE are word overlap-based accuracy metrics that depend on the overall expression of the text, and do not capture the accuracy of the particular key information in the text. Therefore, they may not reflect the factual consistency of a text well, especially for long texts. To alleviate this issue, recent work (Wang et al., 2020; Scialom et al., 2021; Durmus et al., 2020) has proposed new evaluation metrics for abstractive summarization evaluation. These metrics are based on the intuition that if you ask questions about the summary and the original document, you will get a similar answer if the summary realistically matches the original document. They evaluate the accuracy of each local piece of key information in the summary.
@@
We also consider question-answering-based metrics to evaluate the factual accuracy of generation. In this paper, we examine QuestEval (Scialom et al., 2021), a metric that improves the correlation with human judgments over previous metrics in their extensive experiments. QuestEval evaluates the factual consistency between the generated text and the source document, which is mainly used for text summarization tasks. Therefore, it does not require any ground truth reference. However, for RAG systems, the retrieved texts may be irrelevant or incorrect, so consistency with them is not a valid criterion. Instead, we use this metric to measure how well the generated text matches the ground-truth reference. We call this metric RAGQuestEval. We will explain this metric in detail.
@@ 4.1. Experimental Settings
  • Top-k: This is the number of text chunks that the RAG system retrieves for each query, which serves as the context part of the LLMs prompts. The top-k influences the size of the context that the model receives.
@@
Table 3. The experimental results for evaluating different chunk sizes in our benchmark. We use two types of evaluation metrics: The overall semantic similarity metrics(bleu, rouge-L, and bertScore) and the key information metric(RAGQuestEval).
@@ 4.2. Analyzing the Impact of Chunk Size on RAG Performance in Different Tasks
Chunking is the process of dividing a document into chunks of a fixed length, and then converting each chunk into a vector and storing it in an index. This creates an external knowledge index. Chunk size is a crucial parameter that varies depending on the corpus characteristics. If the chunk is too small or too large, it can reduce the search accuracy or omit important content. Hence, finding the optimal chunk size is vital for ensuring search accuracy and relevance, and enabling the LLMs to generate appropriate responses. Our experiments reveal that different RAG tasks correspond to different optimal chunk sizes.
@@
Text Continuation: The experimental results in Table 3 demonstrate that larger chunk size can improve the overall semantic similarity measures (bleu, rouge-L). Besides, the RAGQuestEval metrics, which reflect the precision and recall rate of key information, follow a consistent pattern. This indicates that larger blocks preserve the original document’s structure, which is crucial for creative tasks such as text continuation. Smaller chunks, on the other hand, result in fragmented and semantically incoherent content, which impairs the ability of large models to understand and generate engaging content.
@@
Open-Domain Multi-Document Summarization: We observe some intriguing patterns in the experimental results. Firstly, we discover that larger chunk size not only substantially increases the length of the generated text, but also cause a notable drop in the bleu score, while the rouge-L and bertScore remain almost unchanged. This implies that larger chunks can preserve more original text information, but also introduce some semantic redundancy. Secondly, for the RAGQuestEval metric that evaluates key information, we found that a larger chunk size considerably enhances the recall of key information, but also lowers the precision of key information.
@@
We hypothesize that this is because larger chunks enable the retrieval of more relevant content, thus improving the recall of key information. However, larger chunks also make the summarization task more challenging, as more fine-grained selection is required from the more relevant information, leading to lower precision of key information, which may not be a good thing for the summary.
@@
Question Answering: For single-document QA, too large chunks will reduce both recall and precision score of key information. The task only requires extracting information from a sub-paragraph of a single document, and the answer may be in a specific sentence. Therefore, smaller chunks are more suitable, as excessive content will make the extraction harder for the model.
@@
For multi-document QA, the results are different from those of single-document QA. Larger chunks can significantly improve the recall and precision of key information, as well as the semantic similarity of the generated and reference answers. This is because larger chunks retain the original structure of the article, which is crucial for reasoning and understanding tasks, and fragmented information is not conducive to reasoning.
@@
Hallucination Modification: For the hallucination modification task, the results are similar to those of the single-document QA task. Smaller chunks can significantly improve the semantic similarity metrics, such as the bleu score. This indicates that in the hallucination dataset created by UHGEval, the hallucination information often pertains to only one sentence, which is a mistake at the word or entity level, and does not require the comprehension of long text. Hence, there is no need to understand the whole document, only the relevant portions can be retrieved and modified.
@@ 4.3. Analyzing the Impact of Chunk Overlap on RAG Performance in Different Tasks
Chunk overlap is the number of tokens that two adjacent chunks share. To keep the text semantics coherent, adjacent chunks have some overlap. Chunk overlap determines the size of this overlap. This splitting method meets the maximum length limit of LLMs and maintains the semantic connection between adjacent chunks. Suitable chunk size and overlap can enhance the fluency and coherence of LLMs for long texts. We will show how the chunk overlap rate affects the system performance for different tasks in Table 4.
Text Continuation: With an increase in chunk overlap, we observe a slight enhancement in the metrics that evaluate the alignment of generated text with a reference answer (bleu, rouge-L, and bertScore). The RAGQuestEval metric, which evaluates the accuracy and completeness of important information, improves more obviously. These results indicate that a greater chunk overlap is beneficial for preserving the flow of ideas in the text, which is essential for tasks that require generating new, creative content.
@@
Question Answering: In question answering tasks, chunk overlap has minimal impact on overall semantic similarity metrics such as bleu, rouge-L, and bertScore. However, it significantly affects the accuracy and recall metrics for key information. The results indicate that as chunk overlap increases, the accuracy and recall of key information in single-document question answering tasks improve substantially. Similar improvements are observed in two-document question answering tasks. However, for three-document question answering tasks, the improvement is less pronounced. This may be because three-document question answering tasks require richer context, and larger chunk overlaps may reduce the available context.
@@
Figure 7. Comparison of Mean Reciprocal Rank (MRR) scores for different retrieval methods in our benchmark.